### v2

# TableRAG: A Retrieval Augmented Generation Framework for Heterogeneous Document Reasoning

Abstract
Retrieval-Augmented Generation (RAG) has demonstrated considerable effectiveness in open-domain question answering. However, when applied to heterogeneous documents, comprising both textual and tabular components, existing RAG approaches exhibit critical limitations. The prevailing practice of flattening tables and chunking strategies disrupts the intrinsic tabular structure, leads to information loss, and undermines the reasoning capabilities of LLMs in multi-hop, global queries. To address these challenges, we propose TableRAG, an SQL-based framework that unifies textual understanding and complex manipulations over tabular data. TableRAG iteratively operates in four steps: context-sensitive query decomposition, text retrieval, SQL programming and execution, and compositional intermediate answer generation. We also develop HeteQA, a novel benchmark designed to evaluate the multi-hop heterogeneous reasoning capabilities. Experimental results demonstrate that TableRAG consistently outperforms existing baselines on both public datasets and our HeteQA, establishing a new state-of-the-art for heterogeneous document question answering.
TableRAG: A Retrieval Augmented Generation Framework for Heterogeneous Document Reasoning
Xiaohan Yu∗, Pu Jian∗, Chong Chen#
Huawei Cloud BU, Beijing
{yuxiaohan5, jinapu2, chenchong55}@huawei.com
https://github.com/yxh-y/TableRAG
1 Introduction
Heterogeneous document-based question answering (Chen et al., 2020), which necessitates reasoning over both unstructured text and structured tabular data, presents substantial challenges. Tables are characterized by interdependent rows and columns, while natural language texts are sequential. Bridging this divergence within a unified QA system remains a non-trivial task.
The prevailing approach extends the retrieval-augmented generation (RAG) paradigm, in which the tables are linearized into textual representations (e.g., Markdown) (Gao et al., 2023; Jin and Lu, 2023; Ye et al., 2023b). Typically, chunking strategies are employed (Finardi et al., 2024), wherein flattened tables are segmented and merged with adjacent text spans. During inference, the LLMs generate answers based on the top-N retrieved chunks. However, these methodologies are predominantly tailored to scenarios that require only surface-level comprehension of tables, such as direct answer extraction Pasupat and Liang (2015a); Zhu et al. (2021). When applied to extensive documents that interleave textual and tabular elements, existing RAG methodologies exhibit critical limitations:
Structural Information Loss: The tabular structure integrity is compromised, leading to information loss or irrelevant context that impedes downstream LLMs performance.
Lack of Global View: Due to document fragmentation, the RAG system struggles with multi-hop global queries Edge et al. (2024), such as aggregation, mathematical computations, and other reasoning tasks that require a holistic understanding across entire tables.
As illustrated in Figure 1, the RAG approach computes percentage over the top-N most relevant chunks rather than the full table, and thus results in an incorrect answer.
To address these limitations of existing RAG systems, we propose TableRAG, an SQL-based framework that dynamically transitions between textual understanding and complex manipulations over tabular data. TableRAG interacts with tables by leveraging SQL as an interface. Concretely, the framework operates via a two-stage process: an offline database construction phase and an online inference phase of iterative reasoning. The iterative reasoning procedure comprises four core operations: (i) context-sensitive query decomposition, (ii) text retrieval, (iii) SQL programming and execution, and (iv) intermediate answer generation. The utilization of SQL enables precise symbolic execution by treating table-related queries as indivisible reasoning units, thereby enhancing both computational efficiency and reasoning fidelity. To facilitate rigorous evaluation of multi-hop reasoning over heterogeneous documents, we introduce HeteQA, a novel benchmark consisting of 304 examples across nine diverse domains. Each example contains a composition across five distinct tabular operations. We evaluate TableRAG on both established public benchmarks and our HeteQA dataset against strong baselines, including generic RAG and program-aided approaches. Experimental results demonstrate that TableRAG consistently achieves state-of-the-art performance. Overall, our contributions are summarized as follows:
We identify two key limitations of existing RAG approaches in the context of heterogeneous document question answering: structural information loss and lack of global view.
We propose TableRAG, an SQL-based framework that unifies textual understanding and complex manipulations over tabular data. TableRAG comprises an offline database construction phase and a four-step online iterative reasoning process.
We develop HeteQA, a benchmark for evaluating multi-hop heterogeneous reasoning capabilities. Experimental results show that TableRAG outperforms RAG and programmatic approaches on HeteQA and public benchmarks, establishing a state-of-the-art solution.
2 Task Formulation
In the context of the heterogeneous document question answering task, we define the task input as extensive documents, denoted as where denotes the textual contents and refers to the tabular components. Given a user question , the objective of this task is to optimize a function that, given the combined textual and tabular context, can produce the correct answer :
3 TableRAG Framework
3.1 Overview and Design Principles
We propose TableRAG, an SQL-based framework designed to preserve table structural integrity and facilitate heterogeneous reasoning. As depicted in Figure 2, TableRAG consists of offline and online workflows. The offline phase is tasked with database construction, while the online phase facilitates iterative reasoning. The reasoning procedure unfolds in a four-stage process: (i) Context-sensitive query decomposition, which identifies the respective roles of textual and tabular modalities within the query. (ii) Text retrieval. (iii) SQL programming and execution, which is selectively invoked for subqueries requiring tabular data reasoning. (iv) Compositional intermediate answer generation. The preferential use of SQL is motivated by its capacity to leverage the expressive strength of symbolic execution over structured data, thereby enabling tabular components within user queries to be treated as monolithic reasoning units. In contrast, other languages like Python incur substantial computational overhead when dealing with large-scale data or complex workloads (Shahrokhi et al., 2024).
3.2 Database Construction
In the offline stage, we first extract structured components from heterogeneous documents, yielding a set of tables . To enable information retrieval, we construct two parallel corpora: a textual knowledge base and a tabular schema database. The textual knowledge base comprises both the raw texts and the Markdown-rendered form of each table, denoted as . Both and are segmented into chunks, which are then embedded into dense vector representations using a pre-trained language model (Chen et al., 2024a). For tabular schema database construction, we represent each table by a standardized schema description , derived via a template as follows:
Then, we define a mapping from each flattened table chunk to its originating table schema:
where denotes the -th chunk derived from table . This mapping ensures that local segments remain contextually anchored to the table structure from which they are derived.
The tables are also ingested in a relational database (e.g., MySQL111https://github.com/mysql), supporting symbolic query execution in the subsequent online reasoning.
3.3 Iterative Reasoning
To address multi-hop, global queries that require compositional reasoning over texts and tables, we introduce an iterative inference process aligned with in Equation 1. This process comprises four core operations: (i) context-sensitive query decomposition, (ii) text retrieval, (iii) program and execute SQL, and (iv) compositional intermediate answer generation. Through repeated cycles of decomposition and resolution, a solution to the query is progressively constructed. Detailed prompt templates are provided in Appendix LABEL:prompts.
Context-Sensitive Query Decomposition
We explicitly delineate the respective roles of textual and tabular modalities during the reasoning process. While a table-related query may involve multiple semantic reasoning steps, its tabular resolution can collapse to a single executable operation. Consequently, an effective decomposition of global queries demands more than mere syntactic segmentation, but also structural awareness of the underlying data sources. To this end, we first retrieve the most relevant table content from the textual database and link it to its corresponding table schema description via the mapping function . Based on this, we formulate a subquery at the -th iteration.
Text Retrieval
We deploy a retrieval module that operates in two successive stages: vector-based recall followed by semantic reranking. Given an incoming query , it is encoded into a shared dense embedding space alongside document chunks. We then select the top- candidates with the highest cosine similarity to the query embedding:
In the subsequent reranking stage, the recalled candidate chunks are re-evaluated by a more expressive relevance model, yielding the final top- selections, denoted by .
SQL Programming and Execution
To support accurate reasoning over tabular data, we incorporate a "program-and-execute" mechanism that is selectively invoked only when subquery reasoning involves tables. Specifically, we inspect whether any content originates from tabular sources in the retrieved results. For each chunk in the top-ranked set , we apply the mapping function (in Equation 2) to extract its associated schema, yielding a table schema set:
If the set is empty, this module is passed. Otherwise, we derive an accurate answer with the current subquery and the corresponding schema context as inputs. To achieve this, we leverage structured query execution over relational data and use SQL as the intermediate formal language. A dedicated tool with LLM as backend generates executable SQL programs and applies them to the pre-constructed MySQL database, formalized as follows:
Intermediate Answer Generation
For the subquery , TableRAG can benefit from two heterogeneous information sources: the execution result over SQL database and text retrieval result from the document database. Both of the data sources provide partial or complete evidence. They introduce distinct failure modes: SQL execution may produce incorrect results or execution errors, while text retrieval may yield incomplete or misleading context. Consequently, the results from these sources may either reinforce each other or present contradictions. To address this, we adopt a compositional reasoning mechanism. The execution result and the retrieved textual chunks are cross-examined to validate consistency and guide answer selection. The final answer to each subquery is derived by adaptively weighting the reliability of each source based on its evidential utility, .
Once the query decomposition module determines that no further subqueries are necessary, TableRAG terminates the iterative reasoning process, yielding the final answer , where denotes the total number of iterations performed.
4 Benchmark Construction
In this section, we present HeteQA, a novel benchmark for assessing multi-hop reasoning across heterogeneous documents.
4.1 Data Collection
HeteQA necessitates advanced operations, such as arithmetic computation, nested logic, etc. To balance annotation fidelity with scalability, we adopt a human-in-the-loop collaborative strategy that integrates LLMs with human verification. The construction pipeline proceeds in three stages:
Query Generation
We curate tabular sources from the Wikipedia dataset (Chen et al., 2020). To facilitate analytical depth, we restrict our selection to tables with a minimum of 20 rows and 7 columns, and apply structural deduplication to eliminate redundancy across similar schemas. For each retained table, we define a suite of advanced operations, e.g., conditional filtering, and statistical aggregation. These operations serve as primitives for constructing complex queries. Leveraging the Claude-3.7-sonnet 222https://www.anthropic.com/claude/sonnet, we prompt for query synthesis as compositions over these primitives. Each generated query is paired with executable code in both SQL and Python. We execute the associated code and obtain the answer. A final deduplication pass is applied over both queries and answers, promoting diversity in the dataset. Full implementation details are provided in Appendix A.
Answer Verification
To ensure correctness and reliability, each instance is subjected to manual inspection by human annotators. Their task is to verify that the execution outcome is accurate for the corresponding query. In cases where discrepancies are found, they are responsible for correcting both the underlying code and the resulting answer.
Document Reference
To support queries that integrate both tabular and textual information, we augment the instance by leveraging the associated Wikipedia document. Specifically, certain entities within the query are replaced with reference-based formulations by the human annotators. For example, the query "Which driver " can be rephrased as "What is the nationality of the driver ". This entity substitution can either modify the subject of the question and its corresponding answer or alter the query phrasing while preserving the original answer. The annotation guidelines and annotator profiles are detailed in Appendix LABEL:instructions.
4.2 Discussion
Each data instance in HeteQA is composed of a query, its corresponding answer, the executable SQL sentence, and the execution-derived answer. Through our data collection pipeline, we construct 304 high-quality examples whose answers are grounded in both single-source (82%) and multi-source (18%). The resulting benchmark spans 136 distinct tables and 5314 wiki knowledge entities. To characterize the dataset, we analyze its semantic domains and the types of tabular reasoning operations. As illustrated in Figure 3, HeteQA covers 9 semantically diverse domains and encompasses 5 principal categories of tabular operations. Together, HeteQA constitutes a structurally diverse and semantically broad resource for advancing question answering over heterogeneous documents.
5 Experiments
| Method | Backbone | HybridQA | WikiTQ | HeteQA | ||
| - | - | Single-Source | Multi-Source | Overall | ||
| Direct | Claude-3.5 | 9.84 | 6.21 | 10.68 | 8.65 | 10.00 |
| DeepSeek-R1 | 24.42 | 12.20 | 3.40 | 13.46 | 6.77 | |
| DeepSeek-V3 | 14.75 | 10.39 | 6.80 | 28.85 | 14.19 | |
| Qwen-2.5-72b | 11.47 | 7.37 | 4.85 | 12.50 | 7.42 | |
| NaiveRAG | Claude-3.5 | 20.28 | 82.60 | 33.20 | 40.35 | 34.54 |
| DeepSeek-V3 | 26.56 | 75.40 | 33.60 | 45.61 | 35.85 | |
| Qwen-2.5-72b | 22.62 | 66.33 | 23.07 | 36.84 | 25.66 | |
| ReAct | Claude-3.5 | 43.38 | 69.81 | 26.40 | 44.44 | 29.60 |
| DeepSeek-V3 | 38.36 | 63.40 | 21.14 | 47.39 | 26.07 | |
| Qwen-2.5-72b | 37.38 | 53.80 | 16.94 | 35.71 | 20.47 | |
| TableGPT2 | 9.51 | 63.40 | 35.60 | 16.67 | 32.24 | |
| TableRAG | Claude-3.5 | 47.87 | 84.62 | 44.94 | 40.74 | 44.19 |
| DeepSeek-V3 | 47.87 | 80.40 | 43.32 | 51.85 | 44.85 | |
| Qwen-2.5-72b | 48.52 | 78.00 | 37.65 | 43.96 | 38.82 | |
5.1 Experimental Settings
5.1.1 Datasets.
We assess the performance of TableRAG on our curated HeteQA, as well as multiple established benchmarks spanning two settings:
HybridQA
(Chen et al., 2020) A multi-hop QA dataset involving both tabular and textual information. For our evaluation, we only retain data cases with tables containing more than 100 cells.
WikiTableQuestion
(Pasupat and Liang, 2015b) A TableQA dataset spanning diverse domains. The queries necessitate a range of data manipulation operations, including comparison, aggregation, etc.
5.1.2 Implementation Details
In the text retrieval process, we employ the BGE-M3 series models (Chen et al., 2024a, b). During recall, we retain the top 30 candidates, from which the top 3 are subsequently selected via reranking. To manage large inputs, the text is chunked into segments of 1000 tokens, with a 200-token overlap between consecutive chunks. The iterative loop is bounded by a maximum of 5 iterations. For backbone LLMs, we utilize Claude-3.5-Sonnet as a representative closed-source LLM, while Deepseek-V3, Deepseek-R1 (Guo et al., 2025), and Qwen-2.5-72B (Yang et al., 2024) serve as the open-source counterparts. A consistent backend is maintained for all modules in the online iterative reasoning process. We use accuracy as the evaluation metric, assessed by Qwen-2.5-72B, which yields a binary score of 0 or 1. The prompt is provided in Appendix LABEL:prompts, and the results evaluated using the exact match metric are presented in Appendix LABEL:append_exp.
5.1.3 Baselines
We evaluate the performance of TableRAG by benchmarking it against three distinct baseline methodologies: (1) Direct answer generation with LLMs. (2) NaiveRAG, which processes tabular data as linearized Markdown formatted texts and subsequently applies a standard RAG pipeline. (3) ReAct (Yao et al., 2023), a prompt based paradigm to synergize reasoning and acting in LLMs with external knowledge sources. (4) TableGPT2 (Su et al., 2024) employs a Python-based execution module to generate code (e.g., Pandas) for answer derivation within a simulated environment. The detailed implementation of these baseline methods is provided in Appendix LABEL:baseline and additional baseline comparisons are presented in Appendix LABEL:append_exp.
5.2 Main Result
The main results across different LLMs backbones are presented in Table 1. Several key observations emerge: (1) The ReAct framework demonstrates advantages over naive RAG on multi-source data, but exhibits degraded performance on single-source data that requires tabular reasoning. This can be attributed to context insensitivity during multi-turn reasoning. Queries that could be resolved with single SQL execution are instead decomposed into multiple subqueries. This over-fragmentation may introduce errors or incomplete information during execution (especially when operations such as filtering, aggregation are involved), potentially leading to cascading failures. (2) TableGPT2 yields acceptable results solely on single-source queries, such as WikiTQ, underscoring its limited capacity for handling multi-source queries. This reflects a lack of generalizability in heterogeneous information environments. (3) TableRAG surpasses all baselines, achieving at least a 10% improvement over the strongest alternative. Notably, it performs robustly across both single-source and multi-source data. This performance gain is attributed to the incorporation of symbolic reasoning, which enables effective adaptation to heterogeneous documents. Moreover, the consistency in performance across different LLM backbones underscores TableRAG’s architectural generality and compatibility with a broad range of backbones.
| Method | WikiTQ | HeteQA | ||
| Total Latency | Avg. Latency / Step | Total Latency | Avg. Latency / Step | |
| ReAct | 13.50 | 5.92 | 24.57 | 7.70 |
| TableRAG | 17.70 | 7.93 | 45.65 | 10.79 |
5.3 Ablation Study
To elucidate the relative importance of each component within the TableRAG framework, we evaluate the full architecture against three ablated variants: (1) w/o Context-Sensitive Query Decomposition, where query decomposition is performed without conditioning on retrieved table schema. (2) w/o SQL Execution, which replaces the SQL programming and execution module with the markdown table format. (3) w/o Textual Retrieval, which operates solely through table-based SQL execution, without leveraging textual resources such as Wikipedia documents. The results are summarized in Figure 4. All the modules contribute to the overall performance of TableRAG, though their relative impact varies across benchmarks. On HybridQA, document retrieval proves particularly critical, due to its emphasis on the extraction of entity-centric or numerical cues. Conversely, for HeteQA, SQL execution proves more influential, as the queries involve nested operations that benefit from SQL-based symbolic reasoning. These findings highlight the complementary design of TableRAG’s textual retrieval and program-executed reasoning components.
6 Efficiency
6.1 Execution Dynamics
We evaluate the efficiency of TableRAG by examining the distribution of its execution iterations, as illustrated in Figure 5. Execution lengths are grouped into four categories: fewer than 3 steps, 3–5 steps, exactly 5 steps, and more than 5 steps. Among the evaluated methods, TableGPT2 demonstrates the highest average number of execution steps, with a modal value centered around five. In contrast, TableRAG consistently requires fewer steps, resolving approximately 63.55% of instances in fewer than five steps and an additional 30.00% precisely within five, with only a marginal proportion of cases remaining unsolved under the given iteration constraints. While ReAct exhibits a comparable distribution in execution steps, its overall performance remains markedly inferior to that of TableRAG. These results suggest that TableRAG achieves both superior efficiency in execution and outstanding reasoning accuracy. It is attributed to the incorporation of SQL-based tabular reasoning.
6.2 Latency
We assess the latency performance using Qwen-72B-Instruct as a consistent backbone model. As shown in Table 2, TableRAG incurs lower latency compared to ReAct, in both average time per step and total time efficiency. This is mainly due to: (1) TableRAG executes fewer iterations, and employs symbolic SQL execution, which is computationally efficient. (2) ReAct performs reasoning at every step, and consumes more steps before reaching a final decision, thereby incurring greater time overhead.
7 Analysis
We provide a comprehensive analysis of TableRAG in this section, with additional results presented in Appendix LABEL:analysis.
7.1 Error Analysis
In addition to evaluating the overall performance of TableRAG against established baselines, we performed a detailed error analysis to characterize the nature of prediction failures. Broadly, the incorrect outputs fall into two primary categories: (1) reasoning failures, attributable to errors in SQL execution or flawed intermediate query decomposition, and (2) task incompletion, typically manifesting as refusals to answer or termination upon exceeding the maximum iteration limit. The prediction distribution is shown in Figure 6. Notably, TableGPT2 exhibits the highest frequency of such failures, largely due to its limited capacity to integrate contextual cues from the wiki documents. This constraint frequently results in the model either explicitly refusing to respond or acknowledging its inability to do so. In contrast, ReAct, which lacks mechanisms for context-aware query decomposition and code execution simulation, often engages in unnecessarily elaborate reasoning steps for problems that could be addressed via a single structured inquiry. TableRAG demonstrates the lowest failure rate among the methodologies assessed. Its consistent ability to yield valid responses within five iterations highlights the efficacy of its design — particularly its use of context-aware query decomposition and selective SQL-based execution planning.
7.2 Prediction across Domains
Figure 7 presents a comparative evaluation of TableRAG, instantiated with various backbone LLMs, against the ReAct framework across various domains. The results reveal that TableRAG consistently outperforms ReAct in the majority of domains, demonstrating its effectiveness in heterogeneous document question answering. Only certain domains, such as Culture, exhibit comparatively weaker performance on TableRAG with Qwen backbone. A closer inspection of the data distribution suggests that this degradation may stem from the sparsity of domain-specific instances.
8 Related Work
8.1 Retrieval Augmented Generation
Retrieval-Augmented Generation (RAG) has emerged as a robust paradigm for mitigating hallucination (Zhang et al., 2023a) and enhancing the reliability of Large Language Models (LLMs) generated responses (Lewis et al., 2020; Guu et al., 2020). The RAG approaches retrieve from the knowledge base, and the most relevant document chunks are subsequently incorporated into the generation process (Gao et al., 2023; Zhu et al., 2024; Borgeaud et al., 2022). However, this straightforward retrieval process often yields noisy chunks that may lack critical details, thereby diminishing the quality of the subsequent generation. Recent advancements have thus focused on task-adaptive retrieval mechanisms. Notable frameworks in this regard include Self-RAG (Asai et al., 2023), RQ-RAG (Chan et al., 2024), etc. Despite these innovations, RAG still faces challenges when dealing with heterogeneous contexts (Satpute et al., 2024).
8.2 Table Reasoning via Large Language Models
Table reasoning refers to the development of a system that provides responses to user queries based on tabular data (Lu et al., 2025). The mainstream approaches to table reasoning can be broadly classified into two categories. The first category revolves around leveraging LLMs through prompt engineering. For instance, Tab-CoT (Jin and Lu, 2023) applies chain-of-thought (CoT) reasoning to establish a tabular structured reasoning process. Similarly, Chain-of-Table (Wang et al., 2024) extends the CoT methodology to the tabular setting, enabling a multi-step reasoning process for more complex table-based queries. The second category involves utilizing programs to process tabular data. Tabsqlify (Nahid and Rafiei, 2024) employs a text-to-SQL approach to decompose tables into smaller, contextually relevant sub-tables. DATER (Ye et al., 2023a) adopts a few-shot prompting strategy to reduce large tables into more manageable sub-tables, using a parsing-execution-filling technique that generates intermediate SQL queries. BINDER (Cheng et al., 2022; Zhang et al., 2023b) integrates both Python and SQL code to derive answers from tables. InfiAgent-DABench (Hu et al., 2024) utilizes an LLM-based agent that plans, writes code, interacts with a Python sandbox, and synthesizes results to solve table-based questions.
9 Conclusion
We address the limitations of existing RAG approaches in handling heterogeneous documents that combine textual and tabular data. Current approaches compromise the structural integrity of tables, resulting in information loss and degraded performance in global, multi-hop reasoning tasks. To overcome these issues, we introduce TableRAG, an SQL-driven framework that integrates textual understanding with precise tabular manipulation. To rigorously assess the capabilities of our approach, we also present a new benchmark HeteQA. Experimental evaluations across public datasets and HeteQA reveal that TableRAG significantly outperforms existing baseline approaches.
Limitations
While TableRAG demonstrates strong performance, several limitations merit consideration: 1. The effectiveness of TableRAG is closely tied to the capabilities of the underlying LLMs. Our implementation leverages high-capacity models such as Claude, DeepSeek-v3, and Qwen-72B-Instruct, which possess strong generalization abilities. Smaller models that lack specialized instruction tuning may exhibit a marked degradation in performance. This suggests that achieving competitive results may necessitate substantial computational resources. 2. The HeteQA benchmark is restricted to English. This limitation arises from the difficulty in curating high-quality heterogeneous sources across multiple languages. As a result, cross-lingual generalization remains unexplored. In future work, we aim to extend HeteQA to a multilingual setting, thereby broadening the applicability and robustness of our evaluation framework.
Appendix A HeteQA
| table | List_of_Australian_films_of_2012_0 |
| query | Who wrote and starred the comedy film released in the second half of 2012 (July-December) that had the highest number of cast members in the List of Australian films of 2012? |
| sql_query | Which comedy film released in the second half of 2012 (July-December) had the highest number of cast members in the List of Australian films of 2012? |
| sql |